5) Prepare An Analysis Of Market Strength by calculating for each company the: a) price/earnings ratio b) dividend yield 6) Once you have completed the first 5 steps, write a 1-2 page analysis of the Buckle . What is the strengths, weaknesses, etc.? Why would you invest ot not?

Buckle FEBRUARY 2 2013 (a) JANUARY 31, 2015 FEBRUARY 1, 2014 INCOME STATEMENT DATA $1,153,142 NET SALES $1,128,001 S 1,124,00THE BUCKLE, INC. CONSOLIDATED BALANCE SHEETS (Amounts in Thousands Except Share and Per Share Amounts) January 31 February 1,THE BUCKLE, INC. CONSOLIDATED STATEMENTS OF INCOME (Amounts in Thousands Except Per Share Amounts) Fiscal Years Ended JanuaryB-9 APPENDIX B The Buckle, Inc., 2014 Annual Report THE BUCKLE, INC. CONSOLIDATED STATEMENTS OF COMPREHENSIVE INCOME (AmountsB-10 APPENDIX B The Buckle, Inc., 2014 Annual Report THE BUCKLE, INC. CONSOLIDATED STATEMENTS OF STOCKHOLDERS EQUITY (AmountD. PROPERTY AND EQUIPMENT Shipping/receiving equipment E. FINANCING ARRANGEMENTS 0 6 the amount available for letters of credD. PROPERTY AND EQUIPMENT Shipping/receiving equipment E. FINANCING ARRANGEMENTS 0 6 the amount available for letters of credB-14 APPENDIX B The Buckle, Inc., 2014 Annual Report K. EARNINGS PER SHARE The following table provides reconciliation betwee

Information for #6 :

2) Prepare a Profitability And Total Asset Management Analysis by calculating for each company the: a) profit margin b) asset turnover c) return on assets

A) Profit Margin – Jan 15 = 14.10 Feb 14 = 14.41 Feb 13 = 14.62

B) Asset Turnover – Jan 15 = 2.12 Feb 14 = 2.06 Feb 13 = 2.35

C) Return on Assets – Jan 15 = 29.94 Feb 14 = 29.76 Feb 13 = 34.38

Buckle INC

Jan-15

Feb-14

Feb-13

Net Sales

1153142

1128001

1124007

Net Income

162564

162584

164305

Total Assets

542993

546293

477974

By comparing this ratios we can observe that there was a dip in overall performance from Feb-13 to Feb-14. This have resulted due to increase in assets by 14%. Post this expansion phase their performance is steady in Jan-15.

The net profit of company has remained steady over the observation period meaning thereby they are having steady operations and there expense to income ratio is also stagnant. However it can be observed that increase in assets in Feb-14 has no significant impact on the revenue or the expense. Hence it can be concluded that asset purchase are not used to generate revenue.

Turnover on assets has observed same pattern as return on assets due to steady income to expense ratio.

We can conclude that Buckle Inc. is a safe for investments where investor can expect steady return on their investments.

3) Prepare a Financial Risk Analysis by calculating for each company the: a) debt to equity ratio b) return on equity c) interest coverage ratio*

a) Debt / Equity

Buckle balance sheet debt in the year 2015 = $0 million

Equity in the year = $355.3 million

Debt / Equity in 2015 = 0 / 355.3 = 0%

As company did not have any on balance sheet debt in 2013 and 2014 as well, Debt / Equity ratio was zero in these years as well.

b. Return on Equity = Net Income / Shareholders Equity

For the year 2015, Return on Equity = 162.564 / 355.278 = 45.8%

For the year 2014, Return on Equity = 162.584 / 361.93 = 44.9%

For the year 2013, Return on Equity = 164.305 / 289.649 = 56.7%

c. Interest coverage ratio = Operating profit (EBIT) / Interest expense

As there is no debt, hence no interest expense. Due to this, EBIT is same Income before income tax.For 2015, Interest coverage  = Income before income tax / Interest

4) Prepare a Liquidity Analysis by calculating for each company the cash flow yield a) Cash flows to sales b) Cash flows to assets c) Free cash flows

A) cashflows to sales= operating cash flow/net sales

here free cash flow would be net income.....

so from that data if we calculate cashflow to sales for 2015 would be

for denims would be = 1,62,564/71,040.468

=2.28 times

B) cash flow to assets=cash flow from operstions/average total assets

average total assets= total assets of current year+previous year/2

here weighted average share capital and price of share is given so

total asset of current year (2015)= weighted average shares*price of shares

= 48,090*3.38

= $1,62,544.2

total assets of previous year (2014) = 47,976*3.39 = $1,62,638.64

now average total assets=$1,62,544.2+$1,62,638.64/2 = $1,62,591.42

Cashflow to total assets = 1,62,564/1,62,591.42 = 0.9999 TIMES

IF YOU CANT ANSWER #6 ILL JUST TAKE #5 THANK YOU!!!!!!!

Buckle FEBRUARY 2 2013 (a) JANUARY 31, 2015 FEBRUARY 1, 2014 INCOME STATEMENT DATA $1,153,142 NET SALES $1,128,001 S 1,124,007 S261,699 INCOME BEFORE INCOME TAXES $259,696 $260,456 97,872 PROVISION FOR INCOME TAXES 97,132 97,394 NET INCOME $162,564 $162,584 $ 164,305 DILUTED EARNINGS PER SHARE 3.38 3.39 3.44 NET INCOME AS A PERCENTAGE OF NET SALES 14.6% 14.4% BALANCE SHEET DATA $ 202,318 $218,756 WORKING CAPITAL 147,917 43,698 LONG-TERM INVESTMENTS 43,436 35,735 $546,293 TOTAL ASSETS $542,993 S477,974 LONG-TERM DEBT $ 355,278 $ 361,930 S289,649 STOCKHOLDERS' EQUITY SELECTED OPERATING DATA 460 NUMBER OF STORES OPEN AT YEAR END 450 440 475 AVERAGE SALES PER SQUARE FOOT 459 AVERAGE SALES PER STORE (000'S) 2,321 2,318 2,380 COMPARABLE STORE SALES CHANGE 21% (a) CONSISTS OF 53 WEEKS 2005 $1.13 2005 $501 Z 2006$1.24 2007 O 2007 2008 2009 $620 S1.63 $792 $898 2008 $2.24 $2.73 $2.86 W2009 | 2010 | 2010 $950 U2011 2012 2013 $3.20 S1,124 2012 $3.44 $3.39 2013 Z 2014 $1,128 $1,153

THE BUCKLE, INC. CONSOLIDATED STATEMENTS OF INCOME (Amounts in Thousands Except Per Share Amounts) Fiscal Years Ended January 31, February 1, 2014 February 2, 2013 2015 SALES, Net of returns and allowances of $110,793, $108,851, and $106,612, respectively 1,153,142 S 1,128,001 S,124,007 COST OF SALES (Including buying, distribution, and occupancy costs) 645,810 628,856 624,692 Gross profit 507,332 499,145 499,315 OPERATING EXPENSES: Selling 206,893 201,963 212,688 35,258 39,177 General and administrative 37,671 242,151 241,140 250,359 INCOME FROM OPERATIONS 256,973 256,994 258,175 3,462 OTHER INCOME, Net (Note A) 2,723 3,524 259,696 260,456 INCOME BEFORE INCOME TAXES 261,699 97,132 97,394 PROVISION FOR INCOME TAXES (Note F) 97,872 162,564 S 162,584S 164,305 NET INCOME EARNINGS PER SHARE (Note K): 3.39 S 3.41 S 3.47 asic Diluted 3.38 S 3.39 S 3.44 See notes to consolidated financial statements.
B-9 APPENDIX B The Buckle, Inc., 2014 Annual Report THE BUCKLE, INC. CONSOLIDATED STATEMENTS OF COMPREHENSIVE INCOME (Amounts in Thousands) Fiscal Years Ended January 3 2015 February 1, 2014 February 2, 2013 NET INCOME 162,564 S 162,584 164,305 OTHER COMPREHENSIVE INCOME, NET OF TAX: Change in unrealized loss on investments, net of tax of $240, $56, and S(138), respectively 409 96 (235) Other comprehensive income 409 96 (235) 162.973 S 164,070 COMPREHENSIVE INCOME 162,680 S See notes to consolidated financial statements.
B-10 APPENDIX B The Buckle, Inc., 2014 Annual Report THE BUCKLE, INC. CONSOLIDATED STATEMENTS OF STOCKHOLDERS' EQUITY (Amounts in Thousands Except Share and Per Share Amounts) Accumulated Other Additional Number Common Paid-in Capital Retained of Shares Stock Total la_ Earnings (699) 363,147 BALANCE, January 28, 2012 47,432,089 S 474 100,333 S 263,039 S 164,305 Net income 164,305 Dividends paid on common stock, (SS.30 per share) (254,633) (254,633) 377,520 Common stock issued on exercise of stock options 842 846 Issuance of non-vested stock, net of forfeitures 249,660 Amortization of non-vested stock grants, net of 8,388 8,388 7,831 Income tax benefit related to exercise of stock options 7,831 Change in unrealized loss on investments, net of tax (235) BALANCE, February 2, 2013 48,059,269 S 481 S 17,391 S 72,711 S 4) S 289,649 162,584 Net income 162,584 Dividends paid on common stock, (S2.02 per share) (97,144) (97,144) Common stock issued on exercise of stock options 25,555 Issuance of non-vested stock, net of forfeitures 251,568 Amortization of non-vested stock grants, net of 5,066 5,066 1,679 1,679 Income tax benefit related to exercise of stock options Change in unrealized loss on investments, net of tax 96 96 (838) S 361,930 BALANCE, February 1, 2014 48,336,392 S 483 124,134 S238,151 Net income 162,564 162,564 (176,604) (176,604) Dividends paid on common stock, (S3.66 per share) Common stock issued on exercise of stock options 17,091 70 70 Issuance of non-vested stock, net of forfeitures 26,130 Amortization of non-vested stock grants, net of 6,013 6,013 Income tax benefit related to exercise of stock options 896 896 Change in unrcalized loss on investments, net of tax 409 409 48,379,613 S 484 S 131.112 S 224,111 S (429) S 355,278 BALANCE, January 31,2015 See notes to consolidated financial statements.
D. PROPERTY AND EQUIPMENT Shipping/receiving equipment E. FINANCING ARRANGEMENTS 0 6 the amount available for letters of credit from $20,000 to $10,000 Total future minimum rental commitments under these operating leases with remaining lease terms in excess of one year as of January 31, 2015 are as follows: 0 0 aggregate, to have a material effect on the Company
D. PROPERTY AND EQUIPMENT Shipping/receiving equipment E. FINANCING ARRANGEMENTS 0 6 the amount available for letters of credit from $20,000 to $10,000 Total future minimum rental commitments under these operating leases with remaining lease terms in excess of one year as of January 31, 2015 are as follows: 0 0 aggregate, to have a material effect on the Company
B-14 APPENDIX B The Buckle, Inc., 2014 Annual Report K. EARNINGS PER SHARE The following table provides reconciliation between basic and diluted earnings per share Fiscal Years Ended January 31, 2015 February , 2014 February 2, 2013 Weighted Average Weighted Average Weighted Avcrage Per Share Per Share Per Share Amount Shares- Income Shares AmountInome Amount Shares- Income 47,927 3.39 162,584 Basic EPS S 162,564 47,744 S 3,41 S 164,305 47,383 S 3.47 Effect of Dilutive Securities: Stock options and non- vested shares (0.01) (0.02) 163 232 327 (0.03) Diluted EPS S 162,564 48,090 S 3,38 S 162,584 47976 S 3.39 S 164,305 47,710 S 3.44 No stock options were deemed anti-dilutive and excluded from the computation of diluted camings per share for fiscal 2014, 2013 or 2012 L. SEGMENT INFORMATION The Company is a retailer of medium to better priced casual apparel, footwear, and accessories. The Company operates its business as one reportable segment The Company operated 460 stores located in 44 states throughout the United States as of January 31,2015 The following is information regarding the Company's major product lines, stated as a percentage of the Company's net sales: Fiscal Years Ended January 31 2015 February 1, 2014 February 2 2013 Merchandise Group 43.7% 45.3% 46.4% Denims Tops (including sweaters) 30.8 30.2 30.9 Accessories 8.6 8.4 Fashions 6.2 6.0 5.7 Footwear 2.3 2.3 2.2 Casual bottoms 1.2 0.9 0.8 Other Total 100.0% 100.0% 100.0%
Answer #1

5-

a- price earning ratio =Market price per Share / Earning per Share

= 3.38/3.39=0.997

b- dividend yield ratio= 1- dividend pay out ratio

= 1-dividend pay out ratio

= 1-176604/162564=-0.0863